OpenStudy (anonymous):

Suppose Galileo and his assistant were 1 km apart when Galileo uncovered his lantern to try to measure the speed of light. How long would it have taken the light signal to reach his assistant?

OpenStudy (anonymous):

Light travels at a speed of approximately 3.00 x 10^8 m/s or 3.00 x 10^5 km/s. To find how long it takes for light to travel 1 km through a vacuum, we must divide the numerator and denominator by 3.00 x 10^5 to get 1 km in the numerator and what is in the denominator will be time it takes light to travel that distance in seconds.\[\frac{ \frac{ 3.00 \times 10^5km }{ 3.00 \times 10^5 } }{ \frac{ 1 \sec }{ 3.00 \times 10^5 } } \approx \frac{ 1km }{ 3.33 \times 10^{-6} \sec }\] @Joshuam1

OpenStudy (anonymous):

velocity=distance/time Or, Time=distance/velocity using speed of light 300 000 km/s, \[t=\frac{1}{300,000}=3.3*10^{-6} s\]

OpenStudy (anonymous):

@eseidl Velocity = Displacement / Time Speed = Distance / Time Velocity is a vector, speed is a scalar.

OpenStudy (anonymous):

irrelevant in this problem, but true

OpenStudy (anonymous):

Not completely irrelevant since that information would misguide someone who didn't already know.

OpenStudy (anonymous):

thank you :) both

OpenStudy (anonymous):

very good answers